OpenStudy (keke):

simplify; √(10)*√(20)

OpenStudy (anonymous):

\[10\sqrt{2}\]

OpenStudy (anonymous):

\[\sqrt{10}*\sqrt{20}=\sqrt{10*20}=\sqrt{200}=10*\sqrt{2}\]